OpenStudy (omw2fyb):

Solve the following system of equations by linear combination 3d – e = 7 d + e = 5

KjSaif (kjsaif):

whats the topic? @omw2fyb

KjSaif (kjsaif):

ok :)

KjSaif (kjsaif):

If you add them straight down: 2d-e = 7 d + e = 5 Then the e's cancel out. You are left with 3d = 12 Then solve for d by dividing each side by 3. Now you have d =4. To solve for e go back into one of your original equations and substitute 4 for the d. You can use either equation, so I'll choose the second one. 4 + e = 5 Now solve for e by subtracting each side by 4. Now you have e = 1. So your solutions are d = 4 and e = 1

KjSaif (kjsaif):

~~ http://brainly.com/question/253850